OpenStudy (nuccioreggie):

Compare 6 ⋅ 108 to 3 ⋅ 106. 6 ⋅ 108 is 2,000 times larger than 3 ⋅ 106. 6 ⋅ 108 is 200 times larger than 3 ⋅ 106. 6 ⋅ 108 is 20 times larger than 3 ⋅ 106. 6 ⋅ 108 is 2 times larger than 3 ⋅ 106.

OpenStudy (anonymous):

\[\large 6 \times 108\] or \[\large 6^{108}\]

OpenStudy (nuccioreggie):

I think itts C

OpenStudy (nuccioreggie):

Its right there

OpenStudy (anonymous):

Here, it may be a bit easier if I just wrote them out then \[\large 6 \times 10^{8} = 600,000,000\] \[\large 3 \times 10^{6} = 3,000,000\] \[\large \frac{600,000,000}{3,000,000}=~?\]

OpenStudy (nuccioreggie):

Do i multiply that or what

OpenStudy (anonymous):

Fraction=division

OpenStudy (nuccioreggie):

How do i do it can you write the question out for me but not the answer

OpenStudy (anonymous):

Alternatively, we could have done \[\large \frac{6 \times 10^{8}}{3 \times 10^{6}} \to \frac{6}{3} \times \frac{10^{8}}{10^{6}} \to 2 \times 10^{8-6} \to 2*100 = 200\]

OpenStudy (nuccioreggie):

Ok so its B

OpenStudy (anonymous):

Do you understand WHY though? You've already learned that property of dividing exponents...

OpenStudy (nuccioreggie):

Yes

OpenStudy (nuccioreggie):

Yes